OpenStudy (anonymous):

hi i need help finding the integral of this of 3/x^2+4

OpenStudy (anonymous):

Do you mean\[\int\limits_{}{}\frac{3}{x^2+4}dx\]?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

-3/x+4x +Constant

OpenStudy (anonymous):

Make a substitution of x=2tan(x) and go from there.

OpenStudy (anonymous):

i totally for got can u please walk me through it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Let x=2tan(theta). Then\[\int\limits{}{}\frac{3}{x^2+4}dx=3\int\limits{}{}\frac{d(2\tan \theta)}{4\tan^2 \theta +4}=\frac{3}{4}\int\limits{}{}\frac{2\sec ^2 \theta d \theta}{1+\tan^2 \theta}\]\[=\frac{6}{4}\int\limits{}{}\frac{\sec ^2 \theta d \theta}{\sec^2 \theta}=\frac{3}{2}\int\limits{}{} d \theta= \frac{3}{2} \theta+c\]

OpenStudy (anonymous):

Since\[x = 2\tan \theta \rightarrow \theta = \tan^{-1}\frac{x}{2}\]so\[\int\limits_{}{}\frac{3}{x^2+4}dx= \frac{3}{2}\tan^{-1}\frac{x}{2}+c\]

OpenStudy (anonymous):

ohh the limits of integration are infinity to 1

OpenStudy (anonymous):

Well, you just can use the result ^^. The constant will cancel. In improper integrals, you should set the problem up as\[\lim_{c \rightarrow \infty}\int\limits_{1}^{c}\frac{3}{x^2+4}dx=\lim_{c \rightarrow \infty}\frac{3}{2}\tan^{-1}\frac{x}{2}|_1^c\]

OpenStudy (anonymous):

\[=\lim_{c \rightarrow \infty}\left[ \frac{3}{2}\tan^{-1}\frac{c}{2}-\frac{3}{2}\tan^{-1}\frac{1}{2} \right]\]

OpenStudy (anonymous):

\[=\lim_{c \rightarrow \infty}\frac{3}{2}\tan ^{-1}\frac{c}{2}-\frac{3}{2}\tan^{-1}\frac{1}{2}\]

OpenStudy (anonymous):

\[=\frac{3}{2}.\frac{\pi}{2}-\frac{3}{2}\tan^{-1}\frac{1}{2}=3\pi - \frac{3}{2}\tan^{-1} \frac{1}{2}\]

OpenStudy (anonymous):

hey thank u so much for ur help . can i ask u another quick question

OpenStudy (anonymous):

Wait, the calculation of the limits is wrong.

OpenStudy (anonymous):

hey srry i just checked its 0 to infinity

OpenStudy (anonymous):

so the anser is 3pi/4

OpenStudy (anonymous):

Well, the limit of arctan at infinity is pi/2, and arctan of 0 is 0.

OpenStudy (anonymous):

The answer will be 3pi.

OpenStudy (anonymous):

k ty

OpenStudy (anonymous):

The working for the integral is correct. Change your limits as needed.

OpenStudy (anonymous):

kk great ty can i ask u something very quick about another problem

OpenStudy (anonymous):

if it's quick!

OpenStudy (anonymous):

why does this diverge \[\sum_{2}^{\infty} \] n/n+1

OpenStudy (anonymous):

Since the limit of each term of your sequence does not go to zero, by the nth term test, the series will diverge. If \[\lim_{n \rightarrow \infty}a_n \ne 0\]then\[\sum a_n\] will diverge.

OpenStudy (anonymous):

DON'T CONFUSE THIS WITH THE FOLLOWING: If \[\lim_{n \rightarrow \infty} a_n = 0 \]then\[\sum a_n\]converges.

OpenStudy (anonymous):

im still confused bc if u plug in infinity u get infinity/infinity

OpenStudy (anonymous):

^^ is not always true

OpenStudy (anonymous):

\[\frac{n}{n+1}=\frac{1}{1+1/n}\]

OpenStudy (anonymous):

Now send n to infinity.

OpenStudy (anonymous):

how did u get 1/1+1/n

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\frac{n}{n+1} =\lim_{n \rightarrow \infty}\frac{1}{1+1/n}=\frac{1}{1+0}=1\]

OpenStudy (anonymous):

Divide the numerator AND denominator by n (at the same time). A common practice is to divide the numerator AND denominator with the highest powered term in a polynomial quotient.

OpenStudy (anonymous):

then why is the answer divergent if u got 1

OpenStudy (anonymous):

You can also think of it like this:\[\frac{n}{n+1}=\frac{n}{n(1+1/n)}=\frac{1}{1+1/n}\]since the n's cancel.

OpenStudy (anonymous):

The theorem says, IF you DON'T get 0 when you take the limit of the sequence that makes up the series, then it WON'T converge. 1 is not 0. What it boils down to is a series whose terms aren't dying quick enough for everything to settle down and converge to a limit.

OpenStudy (anonymous):

oh yah i totally forgot

OpenStudy (anonymous):

Don't ever think, though, you can take the limit of a_n and get 0 and your series converges. That's a trap. It ONLY says that if it DOESN'T go to ZERO, then it DEFINITELY WON'T converge.

OpenStudy (anonymous):

the answer is also 1 bc if the n of the top and bottom are to the same power then u can take into account the coefficients

OpenStudy (anonymous):

I'm not too sure what you mean.

OpenStudy (anonymous):

lim as n goes to infinity n/n the answer is 1/1

OpenStudy (anonymous):

I think you need to hone in your algebra skills before moving too much further with sequences and series, since this part of mathematics is all about algebraic manipulation/tricks.

OpenStudy (anonymous):

Yes, since n/n = 1 which is constant.

OpenStudy (anonymous):

But I think you're talking about something called 'asymptotic equivalence'. In the limit, n+1 is approximately equal to n and so it can be expected that, for large n, n/(n+1) is approximately equal to n/n which is 1.

OpenStudy (anonymous):

yah something like that

OpenStudy (anonymous):

The limit as n goes to infinity of 1 is always 1 since it's independent of n.

OpenStudy (anonymous):

Try this site: http://tutorial.math.lamar.edu/Classes/CalcII/SeriesIntro.aspx

OpenStudy (anonymous):

The attachment I sent is on asymptotic equivalence. The method is very helpful when trying to find limits.

OpenStudy (anonymous):

k ty okay im solving another problem and \[\sum_{1}^{\infty}\] 1/2^n+1/n and i don"t know why this is divergent bc when i plug in infi i get 0

OpenStudy (anonymous):

It goes back to what I was saying before - going to 0 is no guarantee of convergence. Only NOT GOING TO ZERO is a guarantee of DIVERGENCE.

OpenStudy (anonymous):

k so i know 1/n does bc the power of n is greater than zero but how do i figure out the integral of 2^n

OpenStudy (anonymous):

isnt the integral of 2^n ln2^n

OpenStudy (anonymous):

srry i meant ln2(2^n)

OpenStudy (anonymous):

You don't need to use integral test. Use limit comparison. Use the harmonic series as your test series to find that\[\lim_{n \rightarrow \infty}\frac{1/n}{1/2^n+1/n}=\lim_{n \rightarrow \infty}\frac{2^n}{2^n+n}=\lim_{n \rightarrow \infty}\frac{2^n}{2^n(1+n/2^n)}\]\[=\lim_{n \rightarrow \infty}\frac{1}{1+n/2^n}=\lim_{n \rightarrow \infty}\frac{1}{1+0}=1\]since this is a non-zero limit, and your test series is a known divergent series, the limit comparison test says that your own series must diverge too.

OpenStudy (anonymous):

alright thank u i got it

OpenStudy (anonymous):

Good :p

OpenStudy (anonymous):

Get your algebra sorted, and use online resources!

OpenStudy (anonymous):

k